### v1

# PRISM: Robust VLM Alignment with Principled Reasoning for Integrated Safety in Multimodality

Abstract
Safeguarding vision-language models (VLMs) is a critical challenge, as existing methods often suffer from over-defense, which harms utility, or rely on shallow alignment, failing to detect complex threats that require deep reasoning. To this end, we introduce PRISM (Principled Reasoning for Integrated Safety in Multimodality), a system2-like framework that aligns VLMs by embedding a structured, safety-aware reasoning process. Our framework consists of two key components: PRISM-CoT, a dataset that teaches safety-aware chain-of-thought reasoning, and PRISM-DPO, generated via Monte Carlo Tree Search (MCTS) to further refine this reasoning through Direct Preference Optimization to help obtain a delicate safety boundary. Comprehensive evaluations demonstrate PRISM’s effectiveness, achieving remarkably low attack success rates including 0.15% on JailbreakV-28K for Qwen2-VL and 90% improvement over the previous best method on VLBreak for LLaVA-1.5. PRISM also exhibits strong robustness against adaptive attacks, significantly increasing computational costs for adversaries, and generalizes effectively to out-of-distribution challenges, reducing attack success rates to just 8.70% on the challenging multi-image MIS benchmark. Remarkably, this robust defense is achieved while preserving, and in some cases enhancing, model utility. To promote reproducibility, we have made our code, data, and model weights available at https://github.com/SaFoLab-WISC/PRISM.
1 Introduction
Figure 1: Helpfulness and harmlessness trade-off on LLaVA-1.5 and Qwen2-VL.
The rapid advancement of Vision-Language Models (VLMs) Liu et al. (2023); Wang et al. (2024b); Zhu et al. (2025) has shown great potential in tasks ranging from image understanding Wang et al. (2024a) and visual question answering Yu et al. (2024); Liu et al. (2024) to complex multimodal reasoning Lu et al. (2024), positioning them as foundational technologies for next-generation AI applications. However, the multimodal nature of these models introduces unprecedented challenges in safety alignment, as malicious actors can exploit vulnerabilities through carefully crafted text, image, or combinations of text and images to lead the model to bypass the internal safety alignment and generate unsafe answers Luo et al. (2024); Wang et al. (2024e; c).
As illustrated in Figure 3, we identify three distinct categories of safety violations in multimodal contexts: (1) Problem unsafe, where the textual prompt contains explicitly harmful content; (2) Image unsafe, where the visual input presents safety risks or violates content policies; and (3) Problem+Image combination unsafe, where neither the text nor image is inherently harmful, but their combination creates safety concerns. This third category represents a particularly challenging threat model, as it exploits the complex semantic relationships between modalities to embed malicious intent within seemingly innocent multimodal inputs, making them invisible to single-modality alignment strategies.
Figure 2: Response comparison between existing defense method with our proposed PRISM method.
Existing defenses typically rely on heuristic safeguards—such as simple keyword filters, visual classifiers, or static refusal rules. These strategies are effective against overtly unsafe text or images, but falter in subtle cross-modal attacks where harm only emerges through the interaction between modalities. The efficacy of training-based defense mechanisms, such as those utilizing Reinforcement Learning from Human Feedback (RLHF) Zhang et al. (2025b); Ji et al. (2025) or Supervised Fine-Tuning (SFT) Zong et al. (2024), is often constrained by their tendency to produce rote refusal responses. A significant limitation of these approaches is their lack of detailed reasoning, which fails to explicitly articulate or reveal the malicious intent underlying a prompt. Consequently, as illustrated in Figure 2, such methods often yield shallow alignment Qi et al. (2025), where they fail to develop a genuine understanding of safety principles and instead rely on superficial pattern matching that can be easily circumvented by sophisticated attacks.
To address this, we take inspiration from System 2-style reasoning that analyzes the interplay between textual and visual elements. Inspired by the success of reasoning to protect large language models Zhang et al. (2025a), effective safety mechanisms should employ chain-of-thought reasoning that carefully examines the context of the problem, analyzes visual content relative to queries, and synthesizes multimodal information to identify safety concerns.
To this end, we introduce PRISM, a safety alignment framework that leverages structured reasoning processes to enhance VLM safety while minimally compromising utility. Our approach consists of two key components: PRISM-CoT, a curated dataset that demonstrates proper safety-aware chain-of-thought reasoning through four structured stages where Problem aims to identify potential harmful content in textual prompts, Caption provides contextualized visual understanding, Reasoning synthesizes multimodal information to detect safety violations, and Output generates appropriate responses with explicit safety justification. PRISM-DPO, a preference optimization dataset generated through Monte Carlo Tree Search (MCTS) that contains step-level preference pairs for direct preference optimization Rafailov et al. (2023).
Our comprehensive evaluations demonstrate that PRISM-DPO significantly enhances the safety of Vision-Language Models (VLMs) while preserving their core utility. The model exhibits robust safety across diverse attack vectors. On static benchmarks like JailbreakV-28K and VLBreakBench, it achieves a near-zero Attack Success Rate (ASR). It also displays strong resilience against adaptive attacks, successfully defending against repeated queries on models like LLaVA-1.5 and thus substantially increasing the attacker’s cost. Furthermore, our framework generalizes effectively to out-of-distribution challenges, achieving a low ASR ($\sim 5\%$) on the multi-image MIS benchmark. Critically, these safety gains do not compromise helpfulness, as our model attains a state-of-the-art score on the MM-Vet-v2 benchmark. As illustrated in Figure 1, our approach strikes an optimal balance between safety and utility, showing that structured reasoning and preference optimization can work synergistically. Finally, our results indicate that safety can be further amplified at test-time by scaling the reasoning framework, highlighting its effectiveness and potential.
2 Related Work
Jailbreaking Vision Language Models. The landscape of Vision-Language Model (VLM) jailbreaking is evolving from simple structural exploits Liu et al. (2024); Luo et al. (2024) to sophisticated attacks that require complex multimodal reasoning. Initial research exposed vulnerabilities through static strategies like Figstep Gong et al. (2025), FC-Attack Zhang et al. (2025c), and MML Wang et al. (2024e), which used crafted input combinations, as well as adaptive frameworks like IDEATOR Wang et al. (2024c). However, a more profound shift is that superficial safety alignments are no longer sufficient, prompting the development of a new generation of benchmarks to probe deeper reasoning vulnerabilities. For instance, MSSBench Zhou et al. (2025) assesses the safety of a language query within its visual context, demanding contextual understanding rather than simple policy adherence. Similarly, VLSBench Hu et al. (2024) requires multimodal cooperation to succeed, focusing on revealing subtle safety cue leakages in text. Furthermore, MIS Ding et al. (2025b) integrates multi-image inputs to test a model’s advanced visual reasoning in complex scenarios. Collectively, these benchmarks signify that the frontier of VLM safety has moved beyond shallow alignment, demanding models with a fundamentally more robust and context-aware reasoning capability.
Safeguarding Vision Language Models. Approaches to safeguarding Vision-Language Models (VLMs) can be broadly categorized into training-based and training-free methods Ding et al. (2025a); Pi et al. (2024). Focusing on training-based approaches, current work primarily follows two paradigms: Supervised Fine-Tuning (SFT) on safety datasets, as seen in Dress Chen et al. (2024b) and VLGuard Zong et al. (2024), and preference optimization using Direct Preference Optimization (DPO), as employed by SPA-VL Zhang et al. (2025b) and SafeRLHF-V Ji et al. (2025). However, by training models for refusal without an explicit reasoning process, these methods often lead to two shortcomings: over-defense, where models incorrectly reject benign queries, and shallow alignment, which leaves them vulnerable to sophisticated attacks that require reasoning to uncover concealed harmful intent.
3 Safety-Aware Reasoning Dataset Generation
In this section, we introduce our reasoning safety dataset generation process, which is designed to create a comprehensive dataset that includes various types of safety violations.
Figure 3: Overview of our reasoning safety dataset generation with three types of safety violations: (1) Problem unsafe where the text prompt contains harmful content, (2) Image unsafe where the visual input presents safety risks, and (3) Problem+Image combination unsafe where the combination of text and image creates safety concerns. […] indicates omitted text for brevity.
As illustrated in Figure 3, we categorize safety violations into three distinct types: (1) Problem unsafe, where the textual prompt contains harmful or inappropriate content, (2) Image unsafe, where the visual input presents safety risks or violates content policies, and (3) Problem+Image combination unsafe, where the interaction between textual and visual modalities creates safety concerns that require sophisticated reasoning to identify the underlying malicious intent.
To be precise, we curate a structured reasoning safety dataset with the format of four CoT reasoning steps: Problem, Caption, Reasoning, and Output.
3.1 Dataset Preparation
Our dataset preparation methodology follows a systematic approach to curating a comprehensive multimodal safety dataset that encompasses the full spectrum of safety violations requiring chain-of-thought reasoning.
Source Dataset Selection. We strategically select three complementary datasets that collectively provide comprehensive coverage of multimodal safety violations: BeaverTails-V Ji et al. (2025), VLGuard Zong et al. (2024), and SPA-VL Zhang et al. (2025b). These datasets were chosen based on their rigorous safety taxonomies, high-quality annotations, and diverse coverage of violation types including harmful textual content, inappropriate visual imagery, and multimodal safety concerns.
Sampling Strategy. To ensure balanced representation across safety categories while maintaining computational feasibility, we employ stratified sampling techniques. From BeaverTails-V, we extract 100 image-text pairs from each violation category within the training partition, ensuring equal representation across all safety dimensions. For VLGuard, we incorporate all training instances labeled as unsafe, yielding approximately 1,000 image-text pairs that provide robust coverage of visual safety violations. From SPA-VL, we randomly sample 1,000 instances from the validation set, leveraging its fine-grained safety category annotations to enhance taxonomic diversity.
Problem-Image Combination Safety Violation Generation. While existing datasets primarily focus on violations at the individual modality level (problem-unsafe or image-unsafe), we recognize a critical gap in addressing combination-unsafe scenarios where neither the text nor image is inherently harmful, but their interaction creates safety concerns requiring sophisticated reasoning. To address this limitation, we adopt the methodology from MML Wang et al. (2024e), which employs steganographic techniques including word replacement, base64 encoding, and rotation to embed malicious content within seemingly benign image-text pairs. This approach creates about 3k instances where the harmful intent is only revealed through careful analysis of the multimodal interaction, necessitating advanced reasoning capabilities for detection.
Through this comprehensive approach, we construct a dataset comprising approximately 7,000 multimodal safety violation instances, categorized into three distinct types: problem-unsafe, image-unsafe, and combination-unsafe scenarios. Then we partition this dataset into 5,000 instances for generating safety-aware chain-of-thought reasoning processes used in supervised fine-tuning, while reserving the remaining as queries for subsequent reinforcement learning phase.
3.2 Safety-aware Chain-of-Thought Reasoning Process Generation
Drawing upon insights from AdaShield Wang et al. (2024d), we recognize the importance of explicitly specifying response methodologies in defense prompts by clearly identifying the specific type of violation that occurs. To this end, we leverage the safety categories annotated in the original datasets and systematically map them to eight distinct safety violation types as defined by Wang et al. (2025). With the given safety categories, we generate a structured chain-of-thought reasoning process that includes four key steps: Problem, Caption, Reasoning, and Output with refusal response once detecting malicious intent, and outputs a refusal that points out the specific type of safety violation.
We employ GPT-4o to generate structured chain-of-thought reasoning processes through a systematic four-step framework:
PROBLEM. This initial step entails a comprehensive examination of the textual prompt to identify potential harmful content or malicious intent. We provide only the text prompt and instruct it to analyze the underlying intention behind the query and determine whether any safety violations are present. Upon detection of a violation, GPT-4o explicitly identifies the specific category of safety infringement according to our established taxonomy.
CAPTION. Recognizing that effective image safety evaluation requires contextual understanding, we instruct GPT-4o to generate a caption that describes the visual content specifically in relation to the problem context. This contextualized captioning approach facilitates the identification of safety violations that might be apparent only when considering the image within the framework of the accompanying query.
REASONING. This critical step involves the integration and synthesis of information derived from both the problem analysis and visual assessment. GPT-4o conducts a systematic evaluation of the interaction between textual and visual modalities, identifying potential safety concerns that emerge specifically from their combination, which is a capability essential for detecting sophisticated multimodal safety violations.
OUTPUT. Upon detection of any safety violation, GPT-4o generates an appropriate refusal response that explicitly articulates the specific type of violation identified in the preceding analytical steps. This ensures alignment with established safety protocols while maintaining transparency regarding the reasoning behind the refusal.
This structured reasoning framework ensures coverage of potential safety violations across individual modalities and their interactions. When malicious intent is detected, subsequent reasoning steps maintain the structural integrity of the process while appropriately refusing to engage with harmful content, thereby creating valuable training examples for our fine-tuning procedure. For a detailed exposition of the prompting methodology, please refer to Appendix B.
Following the generation process, we implement a quality check to verify proper formatting and filter out any jailbroken responses. This yields just under 3,000 high-quality safety reasoning instances. Crucially, this generation process was not unconstrained; we leveraged the carefully crafted safety violation labels from the original datasets as explicit guidance for GPT-4o. By providing these ground-truth labels, we anchor the generated reasoning to human-annotated safety standards, thus ensuring the process is not solely limited by the intrinsic capabilities of the generator model. To preserve performance on benign tasks and prevent over-refusal, we then augment this safety-focused data with 3,000 benign instances from LLaVA-CoT, adapted to our four-step format. This balanced composition results in a principled supervised fine-tuning dataset of approximately 6k instances, which we designate as PRISM-CoT.
Figure 4: Overview of our safety-aware MCTS preference data generation process. (a) Illustrates an image-unsafe instance example, where safety rewards are computed by GPT-4o without back-propagation. (b) Demonstrates a benign instance example, where helpfulness rewards are calculated by GPT-4o using existing reasoning steps as evaluation criteria, with rewards back-propagated through the decision tree.
3.3 Safety-aware MCTS Preference Generation
Monte Carlo Tree Search (MCTS) has been shown to be effective in enhancing LLM’s reasoning capabilities Chen et al. (2024a) and has been successfully applied to enhance LLM’s safety alignment Zhang et al. (2025a). Based on these findings, we further propose to generate Safety-Aware Vision-Language MCTS preference data based on the structured reasoning processes generated in the previous section.
As shown in Figure 4, we generate MCTS preference data for both malicious and benign instances. We construct a reasoning tree where each node $r_{i}^{j}$ represents the $j$-th reasoning step at level $i$, where $i\in\{1,2,3,4\}$ corresponds to Problem, Caption, Reasoning, and Output steps respectively. Each node maintains statistics $(Q(r_{i}^{j}),N(r_{i}^{j}))$ representing the cumulative reward and visit count.
At each node $r_{i}^{j}$, we generate $k$ candidate reasoning steps using our fine-tuned model. Specifically, for a given node $r_{i-1}^{j}$ with partial reasoning sequence $s_{1:i-1}$, we sample $k$ completions for the next reasoning step:
$\{c_{1}^{(i)},c_{2}^{(i)},\ldots,c_{k}^{(i)}\}\sim P_{\theta}(\cdot|s_{1:i-1},\text{image},\text{query})$ (1)
where $P_{\theta}$ represents our PRISM-CoT fine-tuned model and each $c_{\text{1...k}}^{(i)}$ becomes a candidate for $r_{i}^{j}$.
During the selection phase, we choose child nodes using the UCB formula to balance exploration and exploitation:
$\text{UCB}(r_{i}^{j})=\frac{Q(r_{i}^{j})}{N(r_{i}^{j})}+C\sqrt{\frac{\ln N(\text{parent}(r_{i}^{j}))}{N(r_{i}^{j})}}$ (2)
where $C=1.5$ is the exploration constant, $Q(r_{i}^{j})$ is the cumulative reward, $N(r_{i}^{j})$ is the visit count for node $r_{i}^{j}$, and $N(\text{parent}(r_{i}^{j}))$ is the visit count of its parent node.
For safety evaluation, our design philosophy dictates that safety rewards should not undergo back-propagation through the reasoning tree. This approach is motivated by the observation that malicious intent can manifest at various stages of the reasoning process, and penalizing preceding reasoning steps would constitute an inappropriate attribution of blame. Accordingly, we compute the safety reward $R_{s}$ for each node $r_{i}^{j}$ using GPT-4o as an evaluator, which analyzes the reasoning step and assigns a safety score based on the detection of potential malicious intent.
For helpfulness evaluation, we compute the helpfulness reward $R_{h}$ for each node $r_{i}^{j}$ by comparing the current node against the ground truth reference $\text{gt}_{i}^{j}$ derived from established reasoning datasets (i.e., LLaVA-CoT). Unlike safety rewards, helpfulness rewards are propagated back through the reasoning tree via standard MCTS back-propagation mechanisms, enabling iterative refinement of the model’s reasoning capabilities based on response quality assessment.
After multiple MCTS iterations, we collect preference pairs by comparing reasoning paths with different cumulative rewards. For nodes at the same level $i$, we create preference pairs $r_{i}^{j_{1}}\succ r_{i}^{j_{2}}$:
 $\text{ if }Q(r_{i}^{j_{1}})>Q(r_{i}^{j_{2}})+\epsilon\text{ and }Q(r_{i}^{j_{1}})\geq\theta$ (3)
where $\epsilon$ is a small margin to ensure statistical significance and $\theta$ is the threshold for an accepted node.
Following the MCTS preference generation methodology described above, we initiate the process with a carefully curated dataset comprising 500 benign and 1,500 malicious image-text pairs to systematically generate reasoning steps for safety-aware reasoning. Our implementation employs $k=3$ candidate children for each node during the expansion phase, with a maximum iteration limit of $200$ per reasoning tree to ensure computational tractability while maintaining sufficient exploration depth.
For preference pair generation, we establish a difference margin of $\epsilon=0.4$ and a quality threshold of $\theta=0.8$ to identify statistically significant reward differences between reasoning paths. Recognizing that each non-terminal node contributes meaningful insights to the overall reasoning process, we extract preference pairs across all reasoning levels rather than limiting collection to terminal nodes.
Through this systematic preference generation procedure, we obtain a total of 10,000 preference pairs, which we designate as PRISM-DPO. This substantial preference dataset serves as the foundation for our subsequent direct preference optimization training phase.
4 Experiment
4.1 Supervised Fine-tuning on PRISM-CoT
As illustrated in Figure 3, we fine-tune our models on the PRISM-CoT dataset, with each reasoning step delimited by specialized tokens. We augment the model vocabulary with eight special tokens to demarcate the structured reasoning process. The fine-tuning procedure involves comprehensive parameter optimization using $90\%$ of the dataset for training and $10\%$ for validation. We conduct training with a batch size of $16$, a learning rate of 1e51e-5, and train for $5$ epochs. All experiments are performed on $4$ NVIDIA A100 (40GB) GPUs, with each GPU processing $4$ samples per batch.
4.2 Direct Preference Optimization on PRISM-DPO
Following the supervised fine-tuning phase, we proceed with direct preference optimization (DPO) training on the PRISM-DPO dataset. This phase involves optimizing the model to achieve finer-grained alignment at the reasoning step level. For DPO implementation, we employ LoRA fine-tuning with rank $r=16$ and scaling factor $\alpha=64$. To stabilize the training process, we incorporate a mixture of DPO and SFT objectives with a ratio of $1.0$. We utilize a batch size of $16$ and a learning rate of 1e51e-5, conducting training for $3$ epochs across $4$ NVIDIA A100 (40GB) GPUs, with each GPU processing $4$ samples per batch.
4.3 Evaluation
A fundamental challenge in VLM safety is achieving the delicate balance between harmlessness and helpfulness. Models that are overly defensive may reject legitimate queries, diminishing their practical utility, while those prioritizing helpfulness risk responding to harmful requests. Our evaluation specifically addresses this dual objective by assessing both safety robustness against malicious attacks and performance preservation on benign tasks.
Models Methods JailbreakV-28K VLBreak MIS $M$-Vet2
LLM MLLM Challenge Easy Hard
[…]
LLaVA-1.5 No Defense 65.61 22.85 13.00 81.14 78.81 13.1
SafeRLHF-V 64.61 20.00 16.44 74.72 72.48 19.3
SPA-VL 5.64 2.36 6.54 22.84 23.56 20.2
VLGuard 48.82 0.03 1.94 27.85 28.12 12.3
PRISM (Ours) 2.85 0.50 0.20 8.70 20.00 20.4
Qwen2-VL No Defense 15.71 20.23 15.12 71.57 78.02 54.4
SafeRLHF-V 3.13 4.28 10.31 68.56 70.30 12.9
SPA-VL 1.95 5.08 7.82 29.20 30.06 46.8
VLGuard 1.68 0.00 0.04 5.01 7.33 17.7
PRISM (Ours) 1.46 0.07 0.05 3.28 11.29 48.9
Table 1: Results for safety and helpfulness benchmarks. A lower Attack Success Rate signifies greater safety, while a higher GPT-Eval score indicates better helpfulness. Best results are in bold; second-best results are underlined.
Safety Evaluations. To perform a comprehensive safety evaluation, we assess the model across three established benchmarks, each designed to probe distinct multimodal safety vulnerabilities. The first, JailBreakV-28K Luo et al. (2024) is a pioneering benchmark that evaluates both language model-transferred attacks (LLM-Trans) and direct multimodal attacks (MLLM). The second is the Challenge subset of VLBreakBench Wang et al. (2024c), which utilizes an adaptive jailbreaking framework where Gemini-1.5-Pro iteratively generates transferable multimodal attacks against a GPT-4o-mini target, with adversarial images synthesized by Stable Diffusion 3 Esser et al. (2024). The third, the Multi-Image Safety (MIS) benchmark Ding et al. (2025b), introduces a significant challenge by requiring models to reason across multiple images to detect safety violations—a scenario that falls outside the alignment scope of alignment data. We report the Attack Success Rate (ASR), and it was evaluated by llama-guard-3-8B Inan et al. (2023) as the judge for JailbreakV and VLBreak and GPT-4o for MIS by following their settings.
Beyond these static benchmarks, adaptive attacks represent a critical yet often neglected dimension of safety evaluation, as they more accurately reflect real-world scenarios where attackers iteratively refine their strategies based on model responses. Unlike static benchmarks that employ fixed attack patterns, dynamic evaluation frameworks assess model robustness against evolving attack strategies that adaptively exploit model vulnerabilities. To evaluate model robustness against adaptive threats, we employ IDEATOR Wang et al. (2024c) as our dynamic attack framework, which iteratively optimizes adversarial prompts and images. Following the original implementation protocol, we utilize MiniGPT-4 as the attacking agent and evaluate on a harmful corpus: AdvBench Zou et al. (2023). We configure the search parameters with width $N_{w}=7$ and depth $N_{d}=3$, thus having a total of $21$ Attack Queries.
Baselines. We compare our PRISM framework against several key baseline methods for VLM safety. These include SafeRLHF-V Ji et al. (2025), which applies Direct Preference Optimization (DPO) on a dual preference dataset labeled for both helpfulness and safety. Similarly, SPA-VL Zhang et al. (2025b) also leverages DPO, but on a large-scale preference dataset collected from a wide range of models. We also include VLGuard Zong et al. (2024), a method based on Supervised Fine-Tuning (SFT) that uses a dedicated vision-language safety instruction-following dataset.
Helpfulness Evaluation. One critical aspect of safeguarding VLMs is balancing the harmlessness and helpfulness, ensuring that models can effectively assist users while maintaining robust defenses against harmful content. To evaluate this balance, we employ the MM-Vet-v2 benchmark Yu et al. (2024), which provides a comprehensive assessment of multimodal models across various tasks including reasoning, generation, OCR, spatial understanding, knowledge, sequential reasoning, and mathematical capabilities. The benchmark is scored by GPT-4 (gpt-4-0613) with a total score ranging from 0 to 100.
(a) Attack Success Rate (ASR) of an adaptive attack on the AdvBench dataset using MiniGPT-4 as the attacker. The plot tracks the cumulative ASR (y-axis) against the number of queries made to the victim model (x-axis).
(b) Test-time scaling results for Qwen2-VL on the MIS-Challenge subset. The x-axis represents the computational budget multiplier relative to a no-scaling baseline, while the y-axis indicates the achieved safe rate evaluated by GPT-4o.
Figure 5: (a) Adaptive attack robustness and (b) Test-time scaling effectiveness.
Main Results. Our experimental results in Table 1 demonstrate that PRISM achieves a better balance between safety robustness and helpfulness preservation, successfully navigating the critical trade-off in VLM alignment. Across established safety benchmarks, PRISM consistently delivers top-tier performance. On the Qwen2-VL model, for example, it achieves the lowest ASR of $1.46\%$ on JailbreakV-28K (LLM-Trans) and a near-zero ASR on the VLBreak Challenge. While a method like VLGuard also achieves strong safety metrics, it comes at a significant cost to the model’s core utility. This is evident in VLGuard’s drastic drop in helpfulness, where its MM-Vet-v2 score plummets to just $17.7$ for Qwen2-VL, indicating severe over-defense and a loss of general capabilities.
In contrast, PRISM avoids this pitfall. It secures its robust safety profile while maintaining a high level of helpfulness, scoring $20.4$ and $48.9$ on MM-Vet-v2 for LLaVA-1.5 and Qwen2-VL respectively.
Moreover, as illustrated in Figure 5(a), our proposed method demonstrates significant robustness against adaptive attacks. The aligned model consistently maintains a low ASR and requires a large number of queries to find an adversarial example, which translates to a substantially higher computational cost for the attacker. In contrast, the undefended base model is far more vulnerable. Its ASR shows three apparent increases, which correspond to the attack’s architecture: with a search depth of $N_{d}=3$, the attacking agent analyzes past failures to refine its strategy after every $N_{w}=7$ iterations, causing these sharp jumps in ASR.
4.4 Test Time Scaling
The structured nature of the reasoning process enables performance enhancement through test-time scaling. We achieve this by using a self-rewarding method, wherein the model itself functions as a reward model to guide a step-wise scaling process. We then employ Beam Search and Best-of-N strategies to generate and select optimal reasoning paths Snell et al. (2025). The associated computational budget is estimated by the number of candidate comparisons at each reasoning step relative to a baseline without such scaling. By filtering those questions that successfully jailbreak more than half of the defense method in MIS, we get a MIS-Challenge subset to evaluate our PRISM-DPO’s scaling performance on Qwen2-VL.
As illustrated in Figure 5(b), test-time scaling is a highly effective method for improving safety performance. For instance, applying Beam Search with an eightfold (8×) increase in the computational budget enables the model to detect harmful intents that were previously missed by the no-scaling baseline. This enhancement leads to a substantial improvement in the safe rate, which reaches $90\%$ on the MIS-Challenge subset. These findings suggest that allocating greater computational resources during inference is a promising direction for developing more robustly safety-aligned models.
5 Analysis and Ablation
To gain a better understanding of the reason why our PRISM methods are effective, we conduct experiments of analysis and ablation studies.
Models Methods MIS-Hard $M$-Vet2
Qwen2-VL ND 78.02 54.4
PRISM 11.29 48.9
Qwen2.5-VL ND 64.16 59.0
PRISM 6.34 49.2
LLaVA-1.5 ND 78.81 13.1
PRISM 20.00 20.4
LLaVA-1.6 ND 62.97 20.9
PRISM 16.02 27.9
Model JBV-mini MIS
LLM $M$ Easy Hard
Qwen2-VL 14.36 22.36 71.57 78.02
PRISM-DPO 1.54 1.18 3.28 11.29
PRISM-SFT 2.05 1.17 19.33 35.94
 - w/o PU 10.10 3.18 33.12 39.75
 - w/o IU 3.07 13.52 53.10 41.19
 - w/o CU 6.67 9.41 22.13 55.25
Table 2: (a) Effectiveness of our methods across different VLMs, where ND denotes the no-defense baseline method. We report the Attack Success Rate (ASR) on MIS benchmarks and the GPT-Eval score on $M$-Vet-v2. (b) Ablation study examining the contribution of different training data components. We report ASR on JailBreakV-mini and MIS benchmarks, where PU, IU, and CU denote Problem Unsafe, Image Unsafe, and Problem+Image Combination Unsafe data categories.
Our method’s generalizability was tested by extending to more VLMs, including LLaVA-1.6-7B and Qwen2.5-VL-7B. We follow the same training and evaluation settings as described in the previous sections. The results, detailed in Table 2(a), reveal that our PRISM framework exhibits remarkable and consistent effectiveness across different models. Alongside enhancing safety on the most challenging MIS-Hard benchmark with more than 75% ASR decrease, PRISM compromises minimally or even improves utility.
Moreover, to understand the importance of each part in our PRISM framework, we ran an ablation study on the PRISM-CoT dataset. For this test, we trained different versions of our PRISM-SFT model, each time removing one of the three data categories: Problem Unsafe (PU), Image Unsafe (IU), or Problem-Image Combination Unsafe (CU). The results, shown in Table 2(b), clearly indicate that all three parts are necessary to achieve a more comprehensive safety. The most serious drop in safety occurred when we removed the Combination Unsafe (CU) data. This single change caused the Attack Success Rate (ASR) on the MIS-Hard benchmark to jump to $55.25\%$. This result is especially important because the MIS-Hard benchmark is built to test for harmful intent that arises from the connection between the image and the text. Removing the PU and IU data also made the model less safe, confirming that all three data types are critical. This shows they work together to guard the model against many different types of misuse.
6 Conclusion
In this work, we introduce PRISM, a framework designed to protect Vision-Language Models from sophisticated multimodal attacks by safety-aware reasoning. By leveraging such a reasoning process that is refined with direct preference optimization, PRISM achieves state-of-the-art safety across a wide range of benchmarks, including challenging out-of-distribution scenarios. Crucially, these significant safety improvements are realized without degrading the model’s core utility, effectively resolving the critical trade-off between harmlessness and helpfulness. Our findings validate that embedding explicit reasoning is a highly effective and scalable strategy for building VLMs that are both robustly safe and genuinely capable.
Appendix A Case Study
To qualitatively illustrate the performance differences between PRISM and other safety methods, we present a case study below.
The first task involves a benign input requiring both OCR and mathematical reasoning. As shown in Figure 8, the undefended base model attempts the task but arrives at an incorrect answer (20 minutes) due to a minor calculation error, despite a largely correct reasoning process. In contrast, existing defense methods like SPA-VL and VLGuard exhibit significant over-defense; they fail to engage with the reasoning task and instead default to providing only an image caption, highlighting a critical loss of utility. Our PRISM-aligned model shown in Figure 9, however, successfully navigates the complex instructions, providing the correct answer (15 minutes) along with a transparent reasoning process. This case clearly demonstrates PRISM’s ability to maintain essential capabilities on challenging benign tasks, avoiding the utility degradation that affects other safety alignment approaches.
The second case, drawn from the challenging MIS-Hard benchmark, highlights PRISM’s robustness against harmful requests that require contextual reasoning. In this scenario, the SafeRLHF-V model fails, providing a user with detailed instructions on how to graffiti a cultural heritage site—a safety breach (Figure 10). In contrast, our PRISM model successfully identifies the malicious intent during its REASONING stage and delivers a safe refusal, as shown in Figure 11. This case effectively illustrates PRISM’s ability to detect and mitigate complex threats that are embedded in the interaction between text and image, a critical capability where many existing methods prove vulnerable.
Appendix B Dataset Curation Details
In Figure 7, 12, 13, we provide the detailed prompt used in prompting the GPT-4o to generate corresponding reasoning stages with the given labels.
We provide the detailed prompts used to instruct GPT-4o to score the model-generated reasoning steps during the MCTS search. For the helpfulness evaluation, each step is scored against ground truth results from the original dataset using the prompt in Figure 14. The prompts for evaluating the safety score are given in Figures 15 and 16.
Appendix C Helpfulness Evaluation
Table 3 shows results of different defense methods on a benign benchmark, MM-Vet-v2. For models with weaker baselines, such as llava-1.5-7B and llava-1.6-7B, our method PRISM significantly improves benign performance. For instance, on llava-1.5-7B, the total score increases from $13.1$ to $20.4$, and on llava-1.6-7B, it rises from $20.9$ to $27.9$. This suggests that for these models, the benign data included in safety training acts as a beneficial fine-tuning signal, enhancing their general-purpose abilities where they were previously lacking.
In contrast, for highly capable models like Qwen2VL-7B (baseline score of $54.4$), all defenses lead to a performance degradation, highlighting the safety-capability trade-off. Methods like VLGuard and SafeRLHF-V cause a catastrophic performance collapse to $17.7$ and $12.9$, respectively. This is because VLGuard exhibits a high rejection rate for benign queries, while SafeRLHF-V falters on reasoning-intensive tasks like Math. PRISM, however, excels at mitigating this degradation. It maintains a score of $48.9$ on Qwen2VL-7B and $49.2$ on Qwen2.5VL-7B, preserving the models’ original capabilities more effectively than other defenses and thus achieving a better trade-off.
Appendix D Failure Case Analysis
To better understand the limitations of our approach and guide future work, we analyze the failure cases of the Qwen2-VL + PRISM framework on the JailbreakV-28K dataset. As shown in Figure 6, we collected all samples diagnosed as successful jailbreaks by llama-3-guard and visualized their distribution across unsafe categories. The results indicate that jailbreaks related to Non-Violent Crimes and Specialized Advice are the most prevalent. This is a known challenge for safety guardrails, as it is difficult to draw a clear line between harmful content (e.g., advice on financial crimes) and benign queries (e.g., financial advice). Similarly, distinguishing prohibited specialized advice from helpful instructions requires nuanced understanding. Future work should therefore focus on improving the model’s reasoning capabilities to better separate these subtle distinctions.
Models Methods Rec Gen OCR Spat Know Seq Math Total
llava-1.5-7B No Defense 14.9 13.2 8.2 10.7 10.3 4.6 2.9 13.1
SafeRLHF-V 19.2 14.3 14.1 11.9 16.5 6.6 2.9 18.9
SPA-VL 21.9 20.2 11.2 13.7 16.7 11.1 2.6 20.2
VLGuard 13.6 11.6 8.7 9.1 10.5 3.2 5.9 12.3
PRISM (ours) 22.1 18.0 12.7 14.5 18.5 4.1 8.8 20.4
Qwen2VL-7B No Defense 49.4 49.6 62.1 50.5 45.9 35.9 66.5 54.4
SafeRLHF-V 13.5 11.2 9.1 8.7 11.2 10.0 5.9 12.9
SPA-VL 41.6 39.9 52.7 43.2 38.5 30.9 64.4 46.8
VLGuard 13.6 6.5 22.0 23.0 7.3 8.4 46.8 17.7
PRISM (ours) 43.9 41.3 56.6 45.6 40.1 32.7 60.6 48.9
Qwen2.5VL-7B No Defense 54.2 55.2 67.7 55.4 47.8 52.3 69.1 59.0
PRISM (ours) 45.1 46.7 56.2 44.9 42.1 35.3 58.8 49.2
llava-1.6-7B No Defense 20.1 21.7 19.5 14.8 19.9 11.8 17.1 20.9
PRISM (ours) 29.7 28.6 21.7 21.1 26.6 15.2 12.6 27.9
Table 3: Benign performance on MM-Vet-v2 benchmark, scored by GPT-4 (gpt-4-0613).
Figure 6: Distribution of safety violation categories on JailbreakV-28K dataset. Evaluated by llama-3-guard.
Figure 7: Prompt for the definition of safety violation categories.
Figure 8: Example from Qwen2-VL without defense, with SPA-VL and VLGuard on MM-Vet-v2 dataset, with benign input requiring OCR and math capabilities.
Figure 9: Example from Qwen2-VL with our proposed PRISM on MM-Vet-v2 dataset, with benign input requiring OCR and math capabilities.
Figure 10: Example from Qwen2-VL with SafeRLHF-V method on MIS dataset with two images as input.
Figure 11: Example from Qwen2-VL with our proposed PRISM on MIS dataset with two images as input.
Figure 12: Prompt for the generating of reasoning steps (PRISM-CoT).
Figure 13: Prompt for the generating of reasoning steps (PRISM-CoT).
Figure 14: Prompt for scoring helpfulness of the reasoning steps in MCTS process.
Figure 15: Prompt for evaluating safety score of the reasoning steps in MCTS process.
Figure 16: Prompt for evaluating safety score of the reasoning steps in MCTS process.