### v2

# PRISM: Robust VLM Alignment with Principled Reasoning for Integrated Safety in Multimodality

Abstract
Safeguarding vision-language models (VLMs) is a critical challenge, as existing methods often suffer from over-defense, which harms utility, or rely on shallow alignment, failing to detect complex threats that require deep reasoning. To this end, we introduce PRISM (Principled Reasoning for Integrated Safety in Multimodality), a System 2-like framework that aligns VLMs through a structured four-stage reasoning process explicitly designed to handle three distinct categories of multimodal safety violations. Our framework consists of two key components: a structured reasoning pipeline that analyzes each violation category in dedicated stages, and PRISM-DPO, generated via Monte Carlo Tree Search (MCTS) to refine reasoning quality through Direct Preference Optimization. Comprehensive evaluations show that PRISM substantially reduces attack success rates on JailbreakV-28K and VLBreak, improves robustness against adaptive attacks, and generalizes to out-of-distribution multi-image threats, while better preserving model utility on benign multimodal benchmarks. Our code, data, and model weights available at https://github.com/SaFoLab-WISC/PRISM.
PRISM: Robust VLM Alignment with Principled Reasoning
for Integrated Safety in Multimodality
Nanxi Li1, Zhengyue Zhao1, G. Edward Suh2,4, Marco Pavone3,4, Chaowei Xiao1,4 1Johns Hopkins University, 2Cornell University, 3Stanford University, 4NVIDIA
1 Introduction
The rapid progress of Vision-Language Models (VLMs) Liu et al. (2023); Wang et al. (2024b); Zhu et al. (2025) has unlocked strong capabilities across a wide range of tasks, including image understanding Wang et al. (2024a), visual question answering Yu et al. (2024); Liu et al. (2024), and complex multimodal reasoning Lu et al. (2024). These advances position VLMs as foundational technologies for next-generation AI applications. However, their multimodal nature also introduces new and pressing challenges for safety alignment.
Figure 1: Performance of different methods by using LLaVA-1.5 and Qwen2-VL as base models. Our method achieves a better Helpfulness and harmlessness trade-off
While perturbation-based adversarial attacks have been extensively studied Wu et al. (2021); Qi et al. (2024) and can often be mitigated through purification methods Guo et al. (2024) or adversarial training Qi et al. (2024); Cui et al. (2024), structure-based jailbreak attacks Luo et al. (2024); Wang et al. (2025a, 2024c) remain an open challenge. In such attacks, adversaries craft harmful text, images, or text–image combinations to bypass alignment mechanisms and elicit unsafe outputs Luo et al. (2024); Wang et al. (2025a, 2024c). As illustrated in Figure 3, there are three distinct categories of safety violations in multimodal contexts: (1) Problem unsafe, where the textual prompt contains explicitly harmful content; (2) Image unsafe, where the visual input presents safety risks or violates content policies; and (3) Problem+Image combination unsafe, where neither the text nor image is inherently harmful, but their combination creates safety concerns. This third category represents a challenging threat model, as it exploits the complex semantic relationships between modalities to embed malicious intent within seemingly innocent multimodal inputs, making them invisible to single-modality alignment strategies.
Figure 2: Response comparison between the existing defense method with our proposed PRISM method.
Figure 3: Overview of our reasoning safety dataset generation with three types of safety violations: (1) Problem unsafe where the text prompt contains harmful content, (2) Image unsafe where the visual input presents safety risks, and (3) Problem+Image combination unsafe where the combination of text and image creates safety concerns. […] indicates omitted text for brevity.
Existing defenses typically rely on heuristic safeguards—such as simple keyword filters, visual classifiers, or static refusal rules. These strategies are effective against overtly unsafe text or images, but falter in subtle cross-modal attacks where harm only emerges through the interaction between modalities. The efficacy of training-based defense mechanisms, such as those utilizing Reinforcement Learning from Human Feedback (RLHF) Zhang et al. (2025b); Ji et al. (2025) or Supervised Fine-Tuning (SFT) Zong et al. (2024), is often constrained by their tendency to produce rote refusal responses. A significant limitation of these approaches is their lack of detailed reasoning, which fails to explicitly reveal the malicious intent underlying a prompt. Consequently, as illustrated in Figure 2, such methods often yield shallow alignment Qi et al. (2025), where they fail to develop a genuine understanding of safety principles and instead rely on superficial pattern matching that can be easily circumvented by sophisticated attacks.
To address this, we introduce PRISM, a safety alignment framework that moves beyond shallow refusal. The non-triviality of multimodal safety reasoning lies in the fact that neither the text nor the image alone may trigger a safety violation; it is only through their semantic interaction that malicious intent is revealed. PRISM addresses this via two key components:
PRISM-CoT: A curated dataset demonstrating four-stage reasoning (Problem, Caption, Reasoning, Output). This structure is specifically designed to isolate visual context from textual intent, a critical step for detecting combination-unsafe attacks that exploit cross-modal semantic gaps.
PRISM-DPO: A step-level preference optimization dataset generated via Monte Carlo Tree Search (MCTS). This is a principled approach to credit assignment in safety: we initialize the search from the SFT model and apply a stage-local reward strategy. This ensures that a model which reasons correctly about a context is not unfairly penalized for a later refusal or decision error, which is a key driver behind our improved safety–utility trade-off.
Our comprehensive evaluations show that PRISM-DPO substantially improves robustness against multimodal jailbreak attacks while retaining strong performance on benign multimodal tasks. On static benchmarks such as JailbreakV-28K and VLBreakBench, PRISM significantly reduces Attack Success Rate (ASR) compared to the undefended base models and prior training-based defenses. It also improves resilience under adaptive attacks, increasing the number of attacker queries required to find a successful jailbreak. Moreover, PRISM generalizes to out-of-distribution multi-image threats on MIS with consistently lower ASR than competing defenses. Importantly, these safety gains come with markedly better utility retention than existing alignment baselines: across base models, PRISM still maintains good MM-Vet-v2 helpfulness, indicating an improved safety–utility trade-off (Figure 1). Finally, we show safety can be further strengthened via test-time scaling of the structured reasoning framework.
2 Related Work
Jailbreaking Vision Language Models. The landscape of Vision-Language Model (VLM) jailbreaking is evolving from simple structural exploits Luo et al. (2024) to sophisticated attacks that require complex multimodal reasoning. Initial research exposed vulnerabilities through static strategies like Figstep Gong et al. (2025), FC-Attack Zhang et al. (2025c), and MML Wang et al. (2025a), which used crafted input combinations, as well as adaptive frameworks like IDEATOR Wang et al. (2024c). However, a more profound shift is that superficial safety alignments are no longer sufficient, prompting the development of a new generation of benchmarks to probe deeper reasoning vulnerabilities. For instance, MSSBench Zhou et al. (2024) assesses the safety of a language query within its visual context, demanding contextual understanding rather than simple policy adherence. Similarly, VLSBench Hu et al. (2025) requires multimodal cooperation to succeed, focusing on revealing subtle safety cue leakages in text. Furthermore, MIS Ding et al. (2025) integrates multi-image inputs to test a model’s advanced visual reasoning in complex scenarios. Collectively, these benchmarks signify that the frontier of VLM safety has moved beyond shallow alignment, demanding models with a fundamentally more robust and context-aware reasoning capability.
Safeguarding Vision Language Models. Approaches to safeguarding Vision-Language Models (VLMs) can be broadly categorized into training-based and training-free methods Ding et al. (2024); Pi et al. (2024). Focusing on training-based approaches, current work primarily follows two paradigms: Supervised Fine-Tuning on safety datasets, as seen in Dress Chen et al. (2024b) and VLGuard Zong et al. (2024), and preference optimization using Direct Preference Optimization, as employed by SPA-VL Zhang et al. (2025b) and SafeRLHF-V Ji et al. (2025). However, by training models for refusal without an explicit reasoning process, these methods lead to two shortcomings: over-defense, where models incorrectly reject benign queries, and shallow alignment, which leaves them vulnerable to sophisticated attacks that require reasoning to uncover concealed intent.
Reasoning-based Safety Alignment. Recent work demonstrates that embedding explicit reasoning into the alignment process yields more robust safety than shallow refusal training. In the text-only domain, STAIR Zhang et al. (2025a) applies introspective reasoning with MCTS-based preference optimization to improve LLM safety alignment. Deliberative alignment Guan et al. (2024) shows that training models to explicitly reason over safety specifications before responding substantially improves jailbreak robustness while reducing over-refusal. Process-level supervision Lightman et al. (2023) has proven more effective than outcome-only supervision for verifying reasoning chains. Self-rewarding language models Yuan et al. (2024) demonstrate that models can serve as their own reward signals for iterative improvement. While these works establish the value of reasoning for text-only safety, extending this paradigm to the multimodal domain introduces fundamentally distinct challenges: VLMs must jointly reason over visual and textual inputs to detect cross-modal threats where neither modality alone reveals harmful intent. PRISM addresses this gap by designing a multimodal-specific structured reasoning framework with stage-local reward assignment tailored to the unique threat model of vision-language interactions.
3 Safety-Aware Reasoning Dataset Generation
In this section, we introduce our reasoning safety dataset generation process, which is designed to create a comprehensive dataset that includes various types of safety violations.
As illustrated in Figure 3, we categorize safety violations into three distinct types: (1) Problem unsafe, where the textual prompt contains harmful or inappropriate content, (2) Image unsafe, where the visual input presents safety risks or violates content policies, and (3) Problem+Image combination unsafe, where the interaction between textual and visual modalities creates safety concerns that require sophisticated reasoning to identify the underlying malicious intent.
3.1 Dataset Preparation
Our dataset preparation methodology follows a systematic approach to curating a comprehensive multimodal safety dataset that encompasses the full spectrum of safety violations requiring chain-of-thought reasoning. For details about the dataset selection, please refer to the Appendix B.
Problem-Image Combination Safety Violation Generation. While existing datasets primarily focus on violations at the individual modality level (problem-unsafe or image-unsafe), we recognize a critical gap in addressing combination-unsafe scenarios where neither the text nor image is inherently harmful, but their interaction creates safety concerns requiring sophisticated reasoning. To address this limitation, we adopt the methodology from MML Wang et al. (2025a), which employs steganographic techniques including word replacement, base64 encoding, and rotation to embed malicious content within seemingly benign image-text pairs. This approach creates about 3k instances where the harmful intent is only revealed through careful analysis of the multimodal interaction, necessitating advanced reasoning capabilities for detection.
Through this comprehensive approach, we construct a dataset comprising approximately 7,000 multimodal safety violation instances, categorized into three distinct types: problem-unsafe, image-unsafe, and combination-unsafe scenarios. Then we partition this dataset into 5,000 instances for generating safety-aware chain-of-thought reasoning processes used in supervised fine-tuning, while reserving the remaining as queries for subsequent reinforcement learning phase.
3.2 Safety-aware Chain-of-Thought Reasoning Process Generation
Drawing upon insights from AdaShield Wang et al. (2024d), we recognize the importance of explicitly specifying response methodologies in defense prompts by clearly identifying the specific type of violation. We leverage the safety categories annotated in the original datasets and systematically map them to eight distinct safety violation types as defined by Wang et al. (2025b). We then generate structured chain-of-thought reasoning processes through a four-step framework: PROBLEM (analyzing textual intent and identifying safety violations), CAPTION (generating contextualized visual descriptions relative to the query), REASONING (synthesizing cross-modal information to detect multimodal safety concerns), and OUTPUT (producing appropriate refusals with explicit violation justification or helpful answers for benign queries). Detailed descriptions of each step and the prompts are provided in Appendix C.
Crucially, this generation process is grounded in the original datasets rather than relying on the curator model to infer safety intent: for each instance, we provide the human-annotated safety category (and its definition/policy specification) from the source benchmark as explicit conditioning, and require the generated rationale and final decision to be consistent with that label. Following generation, we apply formatting and consistency checks and filter any traces that exhibit unsafe compliance or label-inconsistent decisions, yielding 3,000 high-quality safety reasoning instances. To preserve performance on benign tasks and prevent over-refusal, we augment this safety-focused data with 3,000 benign instances from LLaVA-CoT adapted to our four-step format. This balanced composition results in a principled supervised fine-tuning dataset of approximately 6k instances, which we designate as PRISM-CoT.
Figure 4: Overview of our safety-aware MCTS preference data generation process. (a) Illustrates an image-unsafe instance example, where safety rewards are computed by a judger without back-propagation. (b) Demonstrates a benign instance example, where helpfulness rewards are assigned by a judger using existing reasoning steps as evaluation criteria, with rewards back-propagated through the decision tree.
3.3 Safety-aware MCTS Preference Generation
Monte Carlo Tree Search (MCTS) has been shown to be effective in enhancing LLM’s reasoning capabilities Chen et al. (2024a) and has been successfully applied to enhance LLM’s safety alignment Zhang et al. (2025a). Based on these findings, we further propose to generate Safety-Aware Vision-Language MCTS preference data based on the structured reasoning processes generated in the previous section.
As shown in Figure 4, we generate MCTS preference data for both malicious and benign instances. We construct a reasoning tree where each node $r_{i}^{j}$ represents the $j$-th reasoning step at level $i$, where $i\in\{1,2,3,4\}$ corresponds to Problem, Caption, Reasoning, and Output steps respectively. Each node maintains statistics $(Q(r_{i}^{j}),N(r_{i}^{j}))$ representing the cumulative reward and visit count.
At each node $r_{i}^{j}$, we generate $k$ candidate reasoning steps using our fine-tuned model. Specifically, for a given node $r_{i-1}^{j}$ with partial reasoning sequence $s_{1:i-1}$, we sample $k$ completions for the next reasoning step:
$\{c_{1}^{(i)},c_{2}^{(i)},\ldots,c_{k}^{(i)}\}\sim P_{\theta}(\cdot|s_{1:i-1},\text{image},\text{query})$ (1)
where $P_{\theta}$ represents our PRISM-CoT fine-tuned model and each $c_{\text{1...k}}^{(i)}$ becomes a candidate for $r_{i}^{j}$.
During the selection phase, we choose child nodes using the UCB formula to balance exploration and exploitation:
$\text{UCB}(r_{i}^{j})=\frac{Q(r_{i}^{j})}{N(r_{i}^{j})}+C\sqrt{\frac{\ln N(\text{parent}(r_{i}^{j}))}{N(r_{i}^{j})}}$ (2)
where $C=1.5$ is the exploration constant, $Q(r_{i}^{j})$ is the cumulative reward, $N(r_{i}^{j})$ is the visit count for node $r_{i}^{j}$, and $N(\text{parent}(r_{i}^{j}))$ is the visit count of its parent node.
For safety evaluation, our design philosophy dictates that safety rewards should not undergo back-propagation through the reasoning tree. This approach is motivated by the observation that malicious intent can manifest at various stages of the reasoning process, and penalizing preceding reasoning steps would constitute an inappropriate attribution of blame. Accordingly, we compute the safety reward $R_{s}$ for each node $r_{i}^{j}$ using GPT-4o as an evaluator, which analyzes the reasoning step and assigns a safety score based on the detection of potential malicious intent.
For helpfulness evaluation, we compute the helpfulness reward $R_{h}$ for each node $r_{i}^{j}$ by comparing the current node against the ground truth reference $\text{gt}_{i}^{j}$ derived from established reasoning datasets (i.e., LLaVA-CoT). Unlike safety rewards, helpfulness rewards are propagated back through the reasoning tree via standard MCTS back-propagation mechanisms, enabling iterative refinement of the model’s reasoning capabilities based on response quality assessment.
After multiple MCTS iterations, we collect preference pairs by comparing reasoning paths with different cumulative rewards. For nodes at the same level $i$, we create preference pairs $r_{i}^{j_{1}}\succ r_{i}^{j_{2}}$:
 $\text{ if }Q(r_{i}^{j_{1}})>Q(r_{i}^{j_{2}})+\epsilon\text{ and }Q(r_{i}^{j_{1}})\geq\theta$ (3)
where $\epsilon$ is a margin to ensure statistical significance and $\theta$ is the threshold for an accepted node.
We extract preference pairs across all reasoning levels rather than limiting collection to terminal nodes. Through this procedure, we obtain a total of 10,000 preference pairs, which we designate as PRISM-DPO. Detailed hyperparameters for MCTS generation are provided in Appendix F.
4 Experiment
4.1 Evaluation
Safety Benchmarks. We evaluate on three benchmarks measuring distinct safety dimensions: (1) JailbreakV-28K Luo et al. (2024), covering both LLM-transferred and direct multimodal attacks; (2) VLBreakBench Challenge Wang et al. (2024c), where Gemini-1.5-Pro iteratively generates transferable attacks with adversarial images from Stable Diffusion 3 Esser et al. (2024); and (3) MIS Ding et al. (2025), requiring multi-image reasoning to detect safety violations. ASR is evaluated JailbreakV, VLBreak and MIS, following their original settings. For adaptive attack evaluation, we employ IDEATOR Wang et al. (2024c) with MiniGPT-4 as the attacker on AdvBench Zou et al. (2023).
Baselines. We compare our proposed PRISM, with training details provided in Appendix E, against SafeRLHF-V Ji et al. (2025), which employs Direct Preference Optimization on dual-preference data. Additionally, we evaluate against SPA-VL Zhang et al. (2025b), which utilizes Direct Preference Optimization on preference data. Finally, we include VLGuard Zong et al. (2024), which applies Supervised Fine-Tuning on safety instruction data.
Helpfulness Evaluation. We use MM-Vet-v2 Yu et al. (2024) following their setting to assess utility preservation across reasoning, OCR, spatial understanding, and mathematics tasks.
Models Methods JailbreakV-28K VLBreak MIS $M$-Vet2
LLM MLLM Challenge Easy Hard
[…]
LLaVA-1.5 No Defense 65.61 22.85 13.00 81.14 78.81 13.1
SafeRLHF-V 64.61 20.00 16.44 74.72 72.48 19.3
SPA-VL 5.64 2.36 6.54 22.84 23.56 20.2
VLGuard 48.82 0.03 1.94 27.85 28.12 12.3
PRISM (Ours) 2.85 0.50 0.20 8.70 20.00 20.4
Qwen2-VL No Defense 15.71 20.23 15.12 71.57 78.02 54.4
SafeRLHF-V 3.13 4.28 10.31 68.56 70.30 12.9
SPA-VL 1.95 5.08 7.82 29.20 30.06 46.8
VLGuard 1.68 0.00 0.04 5.01 7.33 17.7
PRISM (Ours) 1.46 0.07 0.05 3.28 11.29 48.9
Table 1: Results for safety and helpfulness benchmarks. A lower Attack Success Rate signifies greater safety, while a higher GPT-Eval score indicates better helpfulness. Best results are in bold; second-best results are underlined.
(a) Attack Success Rate (ASR) of an adaptive attack on the AdvBench dataset using MiniGPT-4 as the attacker. The plot tracks the cumulative ASR (y-axis) against the number of queries made to the victim model (x-axis).
(b) Test-time scaling results for Qwen2-VL on the MIS-Challenge subset. The x-axis represents the computational budget multiplier relative to a no-scaling baseline, while the y-axis indicates the achieved safe rate evaluated by GPT-4o.
Figure 5: (a) Adaptive attack robustness and (b) Test-time scaling effectiveness.
4.2 Main Results.
Table 1 shows that PRISM achieves a favorable trade-off between safety robustness and helpfulness retention. On Qwen2-VL, PRISM reduces JailbreakV-28K (LLM-Trans) ASR to $1.46\%$ and achieves $0.05\%$ ASR on the VLBreak Challenge, while maintaining substantially higher MM-Vet-v2 helpfulness than other safety-aligned baselines (e.g., $48.9$ vs. $46.8$ for SPA-VL and far above $17.7/12.9$ for VLGuard/SafeRLHF-V). On LLaVA-1.5, PRISM achieves consistently low ASR across JailbreakV-28K, VLBreak, and MIS, and attains the highest MM-Vet-v2 score among the evaluated defenses ($20.4$). Overall, while safety alignment can reduce absolute helpfulness for strong base models, PRISM substantially mitigates the utility loss commonly induced by aggressive refusal training, yielding robust safety with competitive benign performance.
Moreover, as illustrated in Figure 5(a), our proposed method demonstrates significant robustness against adaptive attacks. Our models consistently maintain a low ASR and require a large number of queries to find a successful adversarial example, which translates to a substantially higher computational cost for the attacker. In contrast, the undefended base model is far more vulnerable. Its ASR shows three apparent increases, which correspond to the attack’s architecture: with a search depth of $N_{d}=3$, the attacking agent analyzes past failures to refine its strategy after every $N_{w}=7$ iterations, causing these sharp jumps in ASR. However, our method can consistently maintain low ASR.
4.3 Test Time Scaling
The structured nature of the reasoning process enables performance enhancement through test-time scaling. We achieve this by using a self-rewarding method, wherein the model itself functions as a reward model to guide a step-wise scaling process. We then employ Beam Search and Best-of-N strategies to generate and select best reasoning paths Snell et al. (2025). The associated computational budget is estimated by the number of candidate comparisons at each reasoning step relative to a baseline without such scaling. By filtering those questions that successfully jailbreak more than half of the defense method in MIS, we get a MIS-Challenge subset to evaluate our PRISM-DPO’s scaling performance on Qwen2-VL.
As illustrated in Figure 5(b), test-time scaling is a highly effective method for improving safety performance. For instance, applying Beam Search with an eightfold (8×) increase in the computational budget enables the model to detect harmful intents that were previously missed by the no-scaling baseline. This enhancement leads to a substantial improvement in the safe rate, which reaches $90\%$ on the MIS-Challenge subset. These findings suggest that allocating greater computational resources during inference is a promising direction for developing more robustly safety-aligned models.
4.4 Analysis and Ablation
Models Methods MIS-Hard $M$-Vet2
LLaVA-1.6 ND 62.97 20.9
PRISM 16.02 27.9
Qwen2.5-VL ND 64.16 59.0
PRISM 6.34 49.2
Qwen3-VL ND 12.78 66.0
PRISM 2.68 62.5
Model JBV-mini MIS
LLM $M$ Easy Hard
Qwen2-VL 14.36 22.36 71.57 78.02
PRISM 2.05 1.17 19.33 35.94
 - w/o PU 10.10 3.18 33.12 39.75
 - w/o IU 3.07 13.52 53.10 41.19
 - w/o CU 6.67 9.41 22.13 55.25
Table 2: (a) Effectiveness of our methods across different VLMs, where ND denotes the no-defense baseline method. We report the Attack Success Rate (ASR) on MIS benchmarks and the GPT-Eval score on $M$-Vet-v2. (b) Ablation study examining the contribution of different training data components. We report ASR on JailBreakV-mini and MIS benchmarks, where PU, IU, and CU denote Problem Unsafe, Image Unsafe, and Problem+Image Combination Unsafe data categories.
Effectiveness across different base models. We evaluate PRISM on multiple VLM backbones, using the same training protocol and hyperparameters as in prior experiments. As shown in Table 2(a), PRISM consistently reduces the MIS-Hard attack success rate (ASR) while largely preserving helpfulness on MM-Vet2. For example, on LLaVA-1.6, PRISM lowers MIS-Hard ASR from $62.97\%$ to $16.02\%$ and even improves MM-Vet2 from $20.9$ to $27.9$. On Qwen2.5-VL, ASR drops sharply from $64.16\%$ to $6.34\%$ with a moderate MM-Vet2 decrease. Notably, this trend also holds for more recent and already better-aligned backbones: Qwen3-VL starts with a much lower MIS-Hard ASR ($12.78\%$), yet PRISM still achieves a further $4.8\times$ reduction to $2.68\%$ while maintaining high MM-Vet2. Overall, these results suggest PRISM provides transferable safety gains beyond a single base model, including settings where the backbone is strong and comparatively well aligned.
Effectiveness of reasoning step designs. Moreover, to understand the importance of each part in our PRISM framework, we conduct an ablation study on the PRISM-CoT dataset. We trained different versions of our PRISM-SFT model, each time removing one of the three data categories: Problem Unsafe (PU), Image Unsafe (IU), or Problem-Image Combination Unsafe (CU). The results, shown in Table 2(b), clearly indicate that all three parts are necessary to achieve better safety. The most serious drop in safety occurred when we removed the Combination Unsafe (CU) data. This single change caused the Attack Success Rate (ASR) on the MIS-Hard benchmark to jump to $55.25\%$. Removing the PU and IU data also made the model less safe, confirming that all three data types are critical. This shows they work together to align the model against many different types of threats. We further ablate the contribution of each reasoning stage at inference time by appending end tokens to skip specific steps. Results in Appendix Table 4 confirm that every stage contributes: removing Reasoning causes the largest utility drop ($38.21$ vs. $41.27$ on MMMU-Pro), while removing both Problem and Caption causes the largest safety degradation .
Ablating MCTS Structure. Table 3 progressively ablates the key design choices in PRISM-DPO. SFT alone achieves only modest safety gains. Adding naive per-step DPO reduces ASR to $7.8\%$$6.9\%$ but severely degrades utility to $33.0$, falling into the over-defense trap. Introducing MCTS tree structure with safety back-propagation (w SBP) recovers utility to $43.1$, but safety improvement plateaus at $5.1\%$$11.0\%$ ASR—because propagating safety penalties to preceding steps creates incorrect credit assignment: a stage that reasons correctly should not be penalized for a later stage’s safety failure. PRISM-DPO resolves this by keeping safety rewards stage-local (no safety BP) while back-propagating helpfulness rewards, achieving the best of both: lowest ASR with highest utility. This confirms that stage-local safety reward assignment is the critical design principle behind PRISM-DPO’s superior safety–utility balance.
Training JBV-mini MIS $M$-Vet2
Method ASR $\downarrow$ ASR $\downarrow$ Acc $\uparrow$
SFT only 13.2 30.1 38.5
SFT + naive DPO 7.8 6.9 33.0
SFT + MCTS w SBP 5.1 11.0 43.1
SFT + PRISM-DPO 0.38 6.0 48.9
Table 3: Comparison of training configurations on Qwen2-VL. “MCTS w SBP” uses MCTS tree structure with safety reward back-propagation, propagating safety penalties to preceding reasoning steps. PRISM-DPO instead assigns safety rewards stage-locally (no safety BP) and back-propagates only helpfulness rewards, achieving the best safety–utility balance.
5 Conclusion
In this work, we introduce PRISM, a reasoning-based framework designed to protect VLMs from multimodal jailbreak attacks. Across a diverse set of safety benchmarks, PRISM consistently achieves strong robustness and often matches or surpasses prior training-based defenses, including under adaptive and out-of-distribution settings. At the same time, PRISM retains substantially higher benign utility than existing safety alignment baselines, indicating that structured multimodal reasoning can improve the safety–utility trade-off. Our findings suggest that embedding explicit, stage-structured reasoning is an effective and scalable direction for building VLMs that are both safer and practically useful.
Appendix
Appendix A Case Study
To qualitatively illustrate the performance differences between PRISM and other safety methods, we present a case study below.
The first task involves a benign input requiring both OCR and mathematical reasoning. As shown in Figure 8, the undefended base model attempts the task but arrives at an incorrect answer (20 minutes) due to a minor calculation error, despite a largely correct reasoning process. In contrast, existing defense methods like SPA-VL and VLGuard exhibit significant over-defense; they fail to engage with the reasoning task and instead default to providing only an image caption, highlighting a critical loss of utility. Our PRISM-aligned model shown in Figure 9, however, successfully navigates the complex instructions, providing the correct answer (15 minutes) along with a transparent reasoning process. This case clearly demonstrates PRISM’s ability to maintain essential capabilities on challenging benign tasks, avoiding the utility degradation that affects other safety alignment approaches.
The second case, drawn from the challenging MIS-Hard benchmark, highlights PRISM’s robustness against harmful requests that require contextual reasoning. In this scenario, the SafeRLHF-V model fails, providing a user with detailed instructions on how to graffiti a cultural heritage site—a safety breach (Figure 10). In contrast, our PRISM model successfully identifies the malicious intent during its REASONING stage and delivers a safe refusal, as shown in Figure 11. This case effectively illustrates PRISM’s ability to detect and mitigate complex threats that are embedded in the interaction between text and image, a critical capability where many existing methods prove vulnerable.
Appendix B Dataset Curation Details
Source Dataset Selection. We strategically select three complementary datasets that collectively provide comprehensive coverage of multimodal safety violations: BeaverTails-V Ji et al. (2025), VLGuard Zong et al. (2024), and SPA-VL Zhang et al. (2025b). These datasets were chosen based on their rigorous safety taxonomies, high-quality annotations, and diverse coverage of violation types including harmful textual content, inappropriate visual imagery, and multimodal safety concerns.
Sampling Strategy. To ensure balanced representation across safety categories while maintaining computational feasibility, we employ stratified sampling techniques. From BeaverTails-V, we extract 100 image-text pairs from each violation category within the training partition, ensuring equal representation across all safety dimensions. For VLGuard, we incorporate all training instances labeled as unsafe, yielding approximately 1,000 image-text pairs that provide robust coverage of visual safety violations. From SPA-VL, we randomly sample 1,000 instances from the validation set, leveraging its fine-grained safety category annotations to enhance taxonomic diversity.
Appendix C Detailed PRISM-CoT Reasoning Step Descriptions
We employ GPT-4o to generate structured chain-of-thought reasoning processes through the following four-step framework:
PROBLEM. This initial step entails a comprehensive examination of the textual prompt to identify potential harmful content or malicious intent. We provide only the text prompt and instruct the model to analyze the underlying intention behind the query and determine whether any safety violations are present. Upon detection of a violation, the specific category of safety infringement is explicitly identified according to our established taxonomy.
CAPTION. Recognizing that effective image safety evaluation requires contextual understanding, the model generates a caption that describes the visual content specifically in relation to the problem context. This contextualized captioning approach facilitates the identification of safety violations that might be apparent only when considering the image within the framework of the accompanying query.
REASONING. This critical step involves the integration and synthesis of information derived from both the problem analysis and visual assessment. The model conducts a systematic evaluation of the interaction between textual and visual modalities, identifying potential safety concerns that emerge specifically from their combination, which is a capability essential for detecting sophisticated multimodal safety violations.
OUTPUT. Upon detection of any safety violation, an appropriate refusal response is generated that explicitly articulates the specific type of violation identified in the preceding analytical steps. This ensures alignment with established safety protocols while maintaining transparency regarding the reasoning behind the refusal.
This structured reasoning framework ensures coverage of potential safety violations across individual modalities and their interactions. When malicious intent is detected, subsequent reasoning steps maintain the structural integrity of the process while appropriately refusing to engage with harmful content, thereby creating valuable training examples for our fine-tuning procedure.
Curating Prompts. In Figure 7, 12, 13, we provide the detailed prompts used to generate the corresponding reasoning stages with the given labels.
We provide the detailed prompts used to score the model-generated reasoning steps during the MCTS search. For the helpfulness evaluation, each step is scored against ground truth results from the original dataset using the prompt in Figure 14. The prompts for evaluating the safety score are given in Figures 15 and 16.
Appendix D Helpfulness Evaluation
Table 8 shows results of different defense methods on a benign benchmark, MM-Vet-v2. For models with weaker baselines, such as llava-1.5-7B and llava-1.6-7B, our method PRISM significantly improves benign performance. For instance, on llava-1.5-7B, the total score increases from $13.1$ to $20.4$, and on llava-1.6-7B, it rises from $20.9$ to $27.9$. This suggests that for these models, the benign data included in safety training acts as a beneficial fine-tuning signal, enhancing their general-purpose abilities where they were previously lacking.
In contrast, for highly capable models like Qwen2VL-7B (baseline score of $54.4$), all defenses lead to a performance degradation, highlighting the safety-capability trade-off. Methods like VLGuard and SafeRLHF-V cause a catastrophic performance collapse to $17.7$ and $12.9$, respectively. This is because VLGuard exhibits a high rejection rate for benign queries, while SafeRLHF-V falters on reasoning-intensive tasks like Math. PRISM, however, excels at mitigating this degradation. It maintains a score of $48.9$ on Qwen2VL-7B and $49.2$ on Qwen2.5VL-7B, preserving the models’ original capabilities more effectively than other defenses and thus achieving a better trade-off.
Appendix E PRISM Training Details
As illustrated in Figure 3, we first fine-tune our models on the PRISM-CoT dataset for $5$ epochs, with each reasoning step delimited by specialized tokens. We augment the model vocabulary with eight special tokens to demarcate the structured reasoning process. The fine-tuning procedure involves comprehensive parameter optimization using $90\%$ of the dataset for training and $10\%$ for validation. We conduct training with a learning rate of $1\times 10^{-5}$. Following the SFT phase, we proceed with direct preference optimization (DPO) training on the PRISM-DPO dataset. This phase involves optimizing the model to achieve finer-grained alignment at step-level. For DPO implementation, we employ LoRA fine-tuning with rank $r=16$ and scaling factor $\alpha=64$. We utilize a batch size of $16$ and a learning rate of $1\times 10^{-5}$, conducting training for $3$ epochs.
Appendix F MCTS Implementation Details
We initiate the MCTS process with a curated dataset comprising 500 benign and 1,500 malicious image-text pairs. Our implementation employs $k=3$ candidate children for each node during the expansion phase, with a maximum iteration limit of $200$ per reasoning tree to ensure computational tractability while maintaining sufficient exploration. For preference pair generation, we establish a difference margin of $\epsilon=0.4$ and a quality threshold of $\theta=0.8$ to identify statistically significant reward differences between reasoning paths.
Appendix G Ablation on Structured Reasoning Stages
Setting MIS-hard ASR $\downarrow$ MMMU-Pro Acc $\uparrow$
Full PRISM-DPO 11.3 41.27
w/o PROBLEM 19.3 39.42
w/o CAPTION 14.5 40.00
w/o PROBLEM + CAPTION 29.5 36.07
w/o REASONING 20.0 38.21
Table 4: Ablation on structured reasoning stages. Removing any stage degrades both safety and utility, confirming each stage’s contribution.
To isolate whether safety gains stem from the structured reasoning process itself rather than additional supervision, we perform controlled ablations by perturbing each of the four stages (Problem, Caption, Reasoning, Output) via appending end tokens to skip specific steps. As shown in Table 4, the full PRISM pipeline achieves the best balance. Notably, the Reasoning stage is crucial; removing it causes a significant drop in benign utility ($38.21$ vs $41.27$), confirming that improvements reflect genuine reasoning ability rather than just data curation. Ablating any stage increases ASR on MIS-hard, and this is most notable when both Problem and Caption are removed ($29.5$).
Appendix H Comparison with STAIR
Models MMMU-Pro MIS-hard VLBreak
Acc $\uparrow$ ASR $\downarrow$ ASR $\downarrow$
Qwen2-VL (Base) 44.2 78.02 15.12
+ STAIR-1k 41.4 33.19 3.19
+ PRISM-1k 42.9 13.00 0.88
Table 5: Head-to-head comparison between PRISM and STAIR Zhang et al. (2025a) trained on identical 1k data. PRISM’s multimodal-specific design yields substantially lower ASR on cross-modal benchmarks, while also better preserving utility on MMMU-Pro.
STAIR Zhang et al. (2025a) is the closest related work in the text-only safety reasoning literature, following a two-stage SFT+DPO paradigm with MCTS-based preference data generation. While PRISM introduces three principled and complementary contributions that are specifically motivated by the unique challenges of multimodal safety alignment, which STAIR is not designed to address.
Multimodal-specific structured reasoning. STAIR operates on a linear chain-of-thought over textual input only. PRISM instead structures reasoning into four dedicated stages: Problem (analyzing textual intent), Caption (providing a query-aware visual description), Reasoning (cross-modal synthesis to detect interaction-level threats), and Output (generating a justified response). This is a key innovation absent from STAIR: it forces the model to ground its visual understanding relative to the query, which is essential for detecting combination-unsafe scenarios where malicious intent only emerges from the interplay between text and image.
Three-category safety coverage. STAIR’s training data targets text-based safety violations. PRISM explicitly covers three distinct threat categories: Problem-unsafe (harmful text), Image-unsafe (harmful visuals), and Combination-unsafe (benign text and image that become harmful together). This categorical decomposition enables the model to develop targeted reasoning strategies for each threat type, particularly for the most challenging combination-unsafe threats that require sophisticated cross-modal reasoning.
Stage-local safety reward assignment. A critical design choice in PRISM-DPO is that safety rewards are not back-propagated through the MCTS tree. In STAIR, rewards propagate globally across all reasoning steps. This is appropriate for text-only helpfulness objectives, but problematic for safety: a preceding reasoning step that correctly analyzes the context should not be penalized for a later step’s safety failure, as this constitutes incorrect credit assignment. PRISM instead assigns safety rewards locally at each stage, while back-propagating only helpfulness rewards. As demonstrated in Table 3, this design yields the best safety–utility balance.
We conduct a head-to-head comparison by faithfully re-implementing the vanilla STAIR pipeline and training it on the identical 1k data subset used for PRISM. As shown in Table 5, while both methods achieve comparable utility on MMMU-Pro ($41.4$ vs. $42.9$), the safety gap is substantial. STAIR-1k reduces MIS-hard ASR to $33.19\%$, still far above PRISM-1k’s $13.00\%$, a relative improvement of $2.6\times$. On VLBreak, PRISM-1k achieves $0.88\%$ ASR versus $3.19\%$ for STAIR-1k, a $3.6\times$ improvement. The degradation of STAIR on MIS-hard is particularly instructive: this benchmark tests precisely the combination-unsafe threat model where neither text nor image alone reveals the harmful intent. Without a dedicated visual grounding stage, cross-modal threat category coverage, or stage-local reward design, STAIR lacks the mechanisms to reason about such threats, whereas PRISM’s multimodal-specific design directly targets this gap.
Appendix I Evaluator-Agnostic Analysis
PRISM employs GPT-4o in two distinct roles: as the SFT data curator that generates chain-of-thought reasoning traces, and as the reward model that scores candidate reasoning steps during MCTS. Crucially, both roles are grounded to ground truth labels from the original dataset: the SFT annotation is conditioned on human-annotated safety violation categories and their associated safety policy specifications, requiring generated rationales and final decisions to be consistent with those labels; similarly, the MCTS helpfulness judgment scores each candidate step against reference answers derived from the original benchmark. This grounding ensures that neither role relies solely on GPT-4o’s internal safety criteria. We investigate whether these roles nonetheless introduce evaluator-specific artifacts through two controlled experiments targeting each role independently.
Curator-Agnostic SFT Data Generation. We examine whether PRISM-CoT’s reasoning quality is contingent on GPT-4o’s particular reasoning style and safety vocabulary by substituting Qwen2.5-VL, a model from an entirely different organization and training lineage, as the SFT data curator. In both cases, the curator is provided with the same ground truth safety violation categories and policy specifications from the original dataset as explicit conditioning; only the model generating the rationale changes. Qwen2.5-VL generates the structured four-step chain-of-thought traces following the same format specification, while the subsequent MCTS-based DPO stage remains unchanged. As shown in Table 6, replacing the curator yields a final model with MIS-hard ASR of $7.59\%$ versus $6.34\%$ under GPT-4o curation, a marginal degradation of $1.25$ percentage points, while MMMU-Pro accuracy drops by only $0.4$ points ($40.9$ vs. $41.3$). The near-identical post-DPO performance across two curators from different model families demonstrates that PRISM’s safety improvements are driven by the structured reasoning framework itself, rather than by any idiosyncrasy in how GPT-4o articulates safety rationales. It also suggests that the DPO stage is robust to variation in SFT initialization quality.
SFT Data Curator MIS-hard ASR $\downarrow$ MMMU-Pro Acc $\uparrow$
GPT-4o 6.34 41.3
Qwen2.5-VL (Self-Curated) 7.59 40.9
Table 6: Replacing the SFT data curator from GPT-4o with Qwen2.5-VL yields near-identical post-DPO performance (MIS-hard ASR difference of 1.251.25pp; MMMU-Pro difference of 0.40.4pts), confirming that improvement is attributable to the structured reasoning framework rather than to curator-specific artifacts.
Cross-Judge Consistency of MCTS Reward Signals. The MCTS preference pairs by construction encode pairwise ordering judgments: one reasoning path is labeled preferred over another. Importantly, each safety judgment is itself grounded to the human-annotated safety violation category and policy specification from the original dataset, and each helpfulness judgment is scored against the reference answer from the original benchmark, rather than relying on GPT-4o’s unconstrained assessment. If the resulting preference orderings were nonetheless idiosyncratic to GPT-4o’s scoring style, a model trained on these pairs would appear aligned only under GPT-4o evaluation. We directly test this by measuring pairwise preference alignment rates, defined as the fraction of pairs on which two judges agree on the ordering, across four diverse judge models: GPT-4o, GPT-4o-mini, Qwen3-VL, and Gemini-2.5. These judges span three independent organizations (OpenAI, Alibaba, Google), different model scales, and architectures trained on distinct data distributions, making systematic agreement across all pairs unlikely unless the preference signal captures shared, objective safety properties.
As shown in Table 7, all pairwise agreement rates exceed $88.5\%$, with the majority of pairs surpassing $92\%$. Notably, cross-organization pairs such as GPT-4o vs. Qwen3-VL ($92.4\%$) and GPT-4o vs. Gemini-2.5 ($94.0\%$) are comparably high to within-organization pairs. This convergent validity across architecturally and organizationally diverse judges provides strong evidence that the MCTS preference pairs encode reward signals that are grounded in objectively recognizable safety violations rather than in the particular annotation style of any single model. Together, both experiments support the conclusion that PRISM’s safety gains are evaluator-agnostic and structurally robust.
Model GPT-4o 4o-mini Qwen3 Gemini
GPT-4o 100.0 97.1 92.4 94.0
GPT-4o-mini 97.1 100.0 92.0 88.5
Qwen3-VL 92.4 92.0 100.0 90.2
Gemini-2.5 94.0 88.5 90.2 100.0
Table 7: Pairwise preference alignment rates (%) across four judge models drawn from three independent organizations. All rates exceed $88.5\%$, including cross-organization pairs, demonstrating that the MCTS reward signals capture objectively recognizable safety properties rather than GPT-4o-specific annotation artifacts.
Appendix J Failure Case Analysis
To better understand the limitations of our approach and guide future work, we analyze the failure cases of the Qwen2-VL + PRISM framework on the JailbreakV-28K dataset. As shown in Figure 6, we collected all samples diagnosed as successful jailbreaks by llama-3-guard and visualized their distribution across unsafe categories. The results indicate that jailbreaks related to Non-Violent Crimes and Specialized Advice are the most prevalent. This is a known challenge for safety guardrails, as it is difficult to draw a clear line between harmful content (e.g., advice on financial crimes) and benign queries (e.g., financial advice). Similarly, distinguishing prohibited specialized advice from helpful instructions requires nuanced understanding. Future work should therefore focus on improving the model’s reasoning capabilities to better separate these subtle distinctions.
Models Methods Rec Gen OCR Spat Know Seq Math Total
llava-1.5-7B No Defense 14.9 13.2 8.2 10.7 10.3 4.6 2.9 13.1
SafeRLHF-V 19.2 14.3 14.1 11.9 16.5 6.6 2.9 18.9
SPA-VL 21.9 20.2 11.2 13.7 16.7 11.1 2.6 20.2
VLGuard 13.6 11.6 8.7 9.1 10.5 3.2 5.9 12.3
PRISM (ours) 22.1 18.0 12.7 14.5 18.5 4.1 8.8 20.4
Qwen2VL-7B No Defense 49.4 49.6 62.1 50.5 45.9 35.9 66.5 54.4
SafeRLHF-V 13.5 11.2 9.1 8.7 11.2 10.0 5.9 12.9
SPA-VL 41.6 39.9 52.7 43.2 38.5 30.9 64.4 46.8
VLGuard 13.6 6.5 22.0 23.0 7.3 8.4 46.8 17.7
PRISM (ours) 43.9 41.3 56.6 45.6 40.1 32.7 60.6 48.9
Qwen2.5VL-7B No Defense 54.2 55.2 67.7 55.4 47.8 52.3 69.1 59.0
PRISM (ours) 45.1 46.7 56.2 44.9 42.1 35.3 58.8 49.2
llava-1.6-7B No Defense 20.1 21.7 19.5 14.8 19.9 11.8 17.1 20.9
PRISM (ours) 29.7 28.6 21.7 21.1 26.6 15.2 12.6 27.9
Table 8: Benign performance on $M$-Vet-v2 benchmark, scored by GPT-4 (gpt-4-0613).
Figure 6: Distribution of safety violation categories on JailbreakV-28K dataset. Evaluated by llama-3-guard.
Figure 7: Prompt for the definition of safety violation categories.
Figure 8: Example from Qwen2-VL without defense, with SPA-VL and VLGuard on $M$-Vet-v2 dataset, with benign input requiring OCR and math capabilities.
Figure 9: Example from Qwen2-VL with our proposed PRISM on $M$-Vet-v2 dataset, with benign input requiring OCR and math capabilities.
Figure 10: Example from Qwen2-VL with SafeRLHF-V method on MIS dataset with two images as input.
Figure 11: Example from Qwen2-VL with our proposed PRISM on MIS dataset with two images as input.
Figure 12: Prompt for the generating of reasoning steps (PRISM-CoT).
Figure 13: Prompt for the generating of reasoning steps (PRISM-CoT).
Figure 14: Prompt for scoring helpfulness of the reasoning steps in MCTS process.
Figure 15: Prompt for evaluating safety score of the reasoning steps in MCTS process.
Figure 16: Prompt for evaluating safety score of the reasoning steps in MCTS process.